OpenStudy (sophie999nevermind):

The equation of a line is y = –2x + 1. What is the equation of the line that is parallel to the first line and passes through (2, 2)?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (sophie999nevermind):

no...

OpenStudy (skullpatrol):

Please try :-)

OpenStudy (sophie999nevermind):

subtract y from -2x?

OpenStudy (skullpatrol):

What do parallel lines have in common?

OpenStudy (sophie999nevermind):

they both pass 2?

OpenStudy (skullpatrol):

Nope, they have the same slope.

OpenStudy (sophie999nevermind):

whats the final answer

OpenStudy (3mar):

Parallel lines have the same slope and will never intersect.

OpenStudy (3mar):

@sophie999nevermind just walk with skullpatrol step by step to reach your destiny!

OpenStudy (sophie999nevermind):

im guessing its y=1/2x+1

OpenStudy (3mar):

for the same slope: the slope of our line is m=-2 then it passes through the point (2,2) and the formula of a line with a slope m and passes through a point (x_1,y_1) is \[y-y_1=m(x-x_1)\] just plug the coordinates of this point (2,2) in this formula: \[y-2=m(x-2)\] \[y-2=-2(x-2)\] \[y=-2(x-2)+2\] \[y=-2x+4+2\] \[y=-2x+6\]

OpenStudy (skullpatrol):

Your guess would be the equation of a line perpendicular to the given line @sophie999nevermind Next time, making a sketch may help.